Given the demand function PD =27-Q2 and supply function PS =2Q+3 Assuming perfect competition find
1. The comsumers surplus
2. The producer's surplus
1
Expert's answer
2020-10-01T10:00:01-0400

In equilibrium Qd = Qs, so:

27 - 2Q = 2Q + 3,

4Q = 24,

Q = 6 units,

P = 2×6 + 3 = 15.

1. The comsumers surplus is:

"CS = 0.5\u00d7(27 - 15)\u00d76 = 36."

2. The producer's surplus is:

"PS = 0.5\u00d7(15 - 3)\u00d76 = 36."